# Covariant Quantum Error-Correcting Codes with Metrological Entanglement Advantage

### v1

@@ Abstract
We show that a subset of the basis for the irreducible representations of the total $SU(2)$ rotation forms a covariant approximate quantum error-correcting code with transversal $U(1)$ logical gates. Using only properties of the angular momentum algebra, we obtain bounds on the code inaccuracy against generic noise on any known $d$ sites and against heralded $d$-local erasures, generalizing and improving previous works on the “thermodynamic code” to general local spin and different irreducible representations. We demonstrate that this family of codes can host and protect a probe state with quantum Fisher information surpassing the standard quantum limit when the sensing parameter couples to the generator of the $U(1)$ logical gate.
@@
Motivated by these ideas, we introduce a family of AQECCs with transversal continuous gates, based on simple symmetry considerations. Additionally, these codes provide a metrological entanglement advantage for certain code parameters. More specifically, we consider $N$ spin-$s$ degrees of freedom and the irreducible representations (irreps) of their total $SU(2)$ rotation. The irrep basis is $|J,\!M\rangle$, where $J\leq sN$ is the total spin quantum number, and $M$ is the magnetic quantum number. We show that the states $|J,\!M\rangle$ with certain choices of $M$ form an AQECC with transversal $U(1)$ logical gates, providing bounds on its code inaccuracy against generic noise on any known $d$ sites for $J\leq sN$ and against heralded $d$-local erasures for $J=sN$.
Our results stem from generalizing and improving upon results in Refs. [26, 20, 10, 25]. A special case of our AQECCs is $s=\frac{1}{2}$ and $J=sN=\frac{N}{2}$, in which case $|J\!,\!M\rangle$ are the Dicke states [27], and the corresponding code is the “thermodynamic code”. Its code-inaccuracy bound against generic noise on any known $d$ sites was first analyzed in Ref. [26], which also reported a coincidental identical bound for the $(s,J)=(1,N)$ code. We generalize this code to general $s$ and $J\sim N^{a}$ ($a\leq 1$) and tighten the bound against the same noise. We obtain our bound using the $SU(2)$ ladder algebra without explicit expressions for the code words, therefore producing a general $s$- and $J$-dependent bound for this family of codes and explaining the coincidence from Ref. [26].
In addition, consider the state $|J\!=\!\frac{N}{2}\!-\!1,M\!=\!J\rangle=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i2\pi j/N}\hat{q}^{-}_{j}\ket{\uparrow}^{\otimes N}$ along with $|J\!=\!\frac{N}{2}\!-\!1,M\rangle$ for the rest of $M$, where $\hat{q}^{z}_{j}\ket{\uparrow}=\frac{1}{2}\ket{\uparrow}$. In Ref. [28], a subset of the states with $M\sim J-N^{\kappa}$ and $\kappa<1$ was shown to form an approximate error detection code (“magnon code”). Complementing Ref. [28], we consider code words $|J\!=\!\frac{N}{2}\!-\!1,M\rangle$ with $|M|=O(N^{b})$ for some $b<1$.
The inaccuracy bound of the “thermodynamic code” with code words $|M|=O(1)$ in the large-$N$ limit against heralded $d$-local erasure errors for a constant $d$ is reported in Refs. [20, 25]. We generalize it to $s\geq\frac{1}{2}$ and $J=sN$, but with $|M|=O(N^{b})$ and $d\sim N^{c}$ for some exponents $b$ and $c$. We obtain a general $s$-dependent code-inaccuracy bound in this case using the Clebsch-Gordan coefficients without writing out the code words.
@@
Code words from $SU(2)$ irrep states.—Consider a system of $N$ spin-$s$ degrees of freedom (i.e. qudits of dimension $12s\!+\!$ 1), where $s$ can be a half integer or an integer. The total $SU(2)$-rotation unitary is a tensor product of the $SU(2)$ rotations on the individual spins, generated by the $su(2)$ algebra
@@
where $j$ labels the spin and $[\hat{q}_{j}^{+},\hat{q}_{j}^{-}]=2\hat{q}^{z}_{j}$, $[\hat{q}_{j}^{z},\hat{q}_{j}^{\pm}]=\pm\hat{q}_{j}^{\pm}$. This yields $[\hat{Q}^{+},\hat{Q}^{-}]=2\hat{Q}^{z}$ and $[\hat{Q}^{z},\hat{Q}^{\pm}]=\pm\hat{Q}^{\pm}$.
The irreps of the total $SU(2)$-rotation unitary are labeled by $J\leq sN$, with a possible multiplicity, and each irrep has dimension $12J\!+$ 1 with the basis labeled by $|J\!,\!M\rangle$, $M\in\{-J,\cdots,J\}$. The state $|J\!,\!M\rangle$ has the following properties:
@@
We will show that the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, equipped with transversal $U(1)$ logical gates generated by $\hat{Q}^{z}$, forms an AQECC against generic noise on any known $d$ sites and against heralded $d$-local erasures for some parameters. For the maximum $J=sN$, $|J\!,\!M\rangle$ is totally permutation symmetric, and the irrep multiplicity is one. For a general $J<sN$, one can have multiple irreps with the same $J$, and we consider $|J\!,\!M\rangle$ in any irrep or from any linear superposition of different irreps with the same $J$, as long as the ladder algebra in Eq. (2) holds among the resulting $|J\!,\!M\rangle$. As such, we suppress the multiplicity index.
Code inaccuracy against generic errors.—We use the notation $(\!(N,k)\!)$ and $\epsilon(\mathcal{N})$ for an AQECC, where $N$ denotes the number of qudits, $k$ denotes the number of logical qubits, and $\epsilon(\mathcal{N})$ denotes the optimal recovery inaccuracy against noise channel $\mathcal{N}$ (with encoding):
@@
where $D_{P}(\mathcal{E},\mathcal{F})$ is the purified distance between channels $\mathcal{E}$ and $\mathcal{F}$ (see App. S1 for a list of definitions) and $\mathcal{R}$ is the recovery channel. As we only consider noise on at most $d$ sites, we use $\epsilon(d)$ to stress its dependence on $d$, while suppressing $\mathcal{N}$ whose specific choices should be clear from the context.
@@
where $\lambda_{ij}$ are the components of a density operator, and $D_{P}(\Lambda+\mathcal{B},\Lambda)\leq\delta$, where $\Lambda(\rho)=\sum_{i,j}\lambda_{ij}\text{Tr}(\rho)|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\Lambda(\rho)+\sum_{i,j}\text{Tr}(\rho B_{ij})|i\rangle\langle j|$.
@@
Physically, this lemma holds since a $d$-local operator can change the magnetic quantum number $M$ by at most $2sd$. It ensures the “off-diagonal” part of the Knill-Laflamme conditions to be exactly zero if the code words have enough spacing in $M$. The proof is simple and given in App. S2.
@@
Assume $\hat{F}$ is a $d$-local operator, then $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|\leq dq_{0}\|\hat{F}\|_{op}C(n,m)$, where $C(n,m):=\sum_{M=n}^{m-1}[c^{+}_{M}]^{-1}$ and $q_{0}$ is a constant. In particular, for $0\leq b<a\leq 1$, if we take $J\sim N^{a}$ and $|m-n|\sim N^{b}$, then we have $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|=\|\hat{F}\|_{op}O(dN^{b-a})$.
Relegating the proof to App. S3, we comment that this bound is obtained by bounding $|\langle J,n+1|\hat{F}|J,n+1\rangle-\langle J,n|\hat{F}|J,n\rangle|$ and using the triangle inequality. We emphasize that the proof only uses the ladder algebra Eq. (2) and the locality of the operator $Q^{\pm}$ without the explicit expression for $|J\!,\!M\rangle$.
Using these two lemmas, we obtain the following upper bound on the code inaccuracy for generic noise on any known $d$ sites $\mathcal{N}(\rho)=\sum_{j=1}^{n_{d}}K_{j}(\rho)K^{\dagger}_{j}$ whose Kraus operators $K_{j}$ are at most $d$-local, acting on a fixed set of $d$ qudits, and $n_{d}=(2s+1)^{2d}$ is the number of independent Kraus operators. The proof is presented in App. S4.
@@
Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\sim N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, and $b<a\leq 1$. $\mathfrak{C}$ forms an $(\!(N,O(b\log_{2}N))\!)$ AQECC against a $d$-local noise channel with $\epsilon(d)=O(\sqrt{d}(2s+1)^{d}N^{(b-a)/2})$ if we take $\Delta\geq 2sd+1$.
Figure 1: (a) For a $d$-local noise channel, the regimes of $b$ and $c$ with $\epsilon\rightarrow 0$ (blue and red shaded area) and for metrology (red shaded area). (b) For heralded $d$-local erasure errors, the regimes of $b$ and $c$ with $\epsilon\rightarrow 0$ (blue and red shaded area) and for metrology (red shaded area), all independent of $s$.
We focus on parameters where $\epsilon\rightarrow 0$ in the thermodynamic limit $N\rightarrow\infty$. For size $d\sim c\log_{2}N$, we have $\epsilon=O(\sqrt{\log_{2}N}N^{(2c\log_{2}(2s+1)+b-a)/2})$, requiring $c<(a-b)/(4\log_{2}(2s+1))$ and $b<a$ to have $\epsilon\rightarrow 0$ when $N\rightarrow\infty$ as shown in Fig. 1(a) as the shaded area.
@@
The reason to consider the noise on a fixed set of $d$ sites is to control the number $n_{d}$. For generic $d$-local noise $\mathcal{N}(\rho)=\sum_{j=1}^{n_{d}}K_{j}(\rho)K^{\dagger}_{j}$, where $K_{i}$ can act on different sets of $d$ qudits, $\mathfrak{C}$ forms an $((N,O(b\log_{2}N)))$ AQECC with $\epsilon(d)=O(\sqrt{d}n_{d}N^{(b-a)/2})$ if we take $\Delta\geq 4sd+1$. (See App. S4 for this corollary and its proof.)
Code inaccuracy against heralded erasures.—The inaccuracy of the “thermodynamic code” against heralded $d$-local erasures for $d=O(1)$ and $|M|=O(1)$ was found in Refs. [20, 25] to be $\epsilon=O(N^{-1})$. The heralded $d$-local erasure channel is $\mathcal{N}_{\alpha}(\rho):=\text{Tr}_{\alpha}[\rho]$, where $\alpha$ denotes the set of $d$ erased qudits.
We now generalize the code-inaccuracy bound against heralded erasures to $s$, $d\sim N^{c}$ and $|M|=O(N^{b})$. In this case, we consider codes with $J=sN$ only, where $|J\!,\!M\rangle$ is totally permutation symmetric. To bound the code inaccuracy, we follow Ref. [18] by bounding the channel distance via the complementary channel $\hat{\mathcal{N}}_{\alpha}(\bullet)=\text{Tr}_{\bar{\alpha}}(\bullet)$, where $\bar{\alpha}$ is the complement of $\alpha$. This can be achieved by the following lemma.
@@
Consider the irrep of $SU(2)$ formed by $|J\!,\!M\rangle$, where $J=sN$, and the reduced density matrix $\rho_{M}:=\text{Tr}_{\bar{\alpha}}[|J\!,\!M\rangle\langle J\!,\!M|]$, where $\bar{\alpha}$ is the complement of the set of erased qudits $\alpha$. We then have the fidelity $f(\rho_{M},\rho_{M=0})=1-O(\frac{dM^{2}}{sN^{2}})=1-O(N^{2b+c-2})$ asymptotically for $M\sim N^{b}$, $d\sim N^{c}$, and $1>b\geq c$.
Physically, the above lemma states that the environment can barely tell the code words apart from each other. In the proof (presented in App. S5), we use Clebsch-Gordan coefficients without writing the code words out explicitly, enabling a general $s$-dependent bound on the fidelity for $J=sN$. In particular, we can rewrite $|J\!,\!M\rangle=\sum_{m_{1}=-j_{1}}^{j_{1}}C^{JM}_{j_{1},m_{1},j_{2},m_{2}}|j_{1}\!,\!m_{1}\rangle_{\alpha}\otimes|j_{2}\!,\!m_{2}\rangle_{\bar{\alpha}}$, where $m_{2}=M-m_{1}$, $j_{1}=sd$, $j_{2}=J-j_{1}$, and $C^{JM}_{j_{1},m_{1},j_{2},m_{2}}=(\langle j_{1},m_{1}|\langle j_{2},m_{2}|)|J,M\rangle$ is the Clebsch-Gordan coefficient. The fidelity between $\rho_{M}$ and $\rho_{M=0}$ can therefore be calculated accordingly.
@@
Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J=sN$, $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $b<1$. $\mathfrak{C}$ forms an $(\!(N,O((b\!-\!c)\log_{2}N))\!)$ AQECC against heralded $d$-local erasures of size $d\sim N^{c}$, where $10\leq c\leq b<$ 1, with $\epsilon(d)=O(N^{b+c/2-1})$ if we take $\Delta\geq 2sd+1$.
Theorem 2 can be obtained by combining our Lemma 1, our Lemma 3, and Theorem 3 of Ref. [20]. To make our presentation self-contained, we present the proof of Theorem 2 tailored to our case in App. S6. Refs. [20, 25] showed that $\epsilon=\Omega(N^{-1})$ for any $U(1)$ covariant code. Thus, this family of $SU(2)$ AQECCs is asymptotically optimal for general $s$ since $\epsilon=O(N^{-1})$ when $b=c=0$.
For general $b$ and $c$, to have $\epsilon\rightarrow 0$ when $N\rightarrow\infty$, we need $c<2(1-b)$. Additionally, we need $\Delta=\Omega(N^{c})$ for Lemma 1 and $c\leq b$ to ensure at least one logical qubit. This parameter regime is shaded in Fig. 1(b), which is independent of $s$ and $k$. The code words with $M\sim N^{b}$ for $b>\frac{1}{2}$ [red shaded area in Fig. 1(b)] are crucial for a metrological entanglement advantage.
Quantum metrological advantage.—We now apply our codes to quantum metrology and show that they can have an entanglement advantage with noise resistance. In the local parameter estimation problem [10], we would like to estimate the parameter $\theta$ imprinted on the probe state $|\psi_{0}\rangle\in\mathfrak{C}$ as $|\psi_{\theta}\rangle=e^{-i\hat{Q}^{z}\theta}|\psi_{0}\rangle\in\mathfrak{C}$. Abbreviating $\psi_{\theta}:=|\psi_{\theta}\rangle\langle\psi_{\theta}|$, the noisy probe state is then $\rho_{\theta}:=\mathcal{N}(\psi_{\theta})$ for the noise channel $\mathcal{N}$.
In quantum metrology, QFI is a central figure of merit, which characterizes how well a parameter can be estimated via the quantum Cramér-Rao bound [41]. For the states $\rho_{\theta}$ parameterized by $\theta$, the QFI is defined as
$\mathfrak{F}(\rho_{\theta};\partial_{\theta}\rho_{\theta}):=\text{Tr}(\rho_{\theta}R^{2})~{},$ (5)

### v2

@@ Abstract
We show that a subset of the basis for the irreducible representations of a tensor-product SU(2) rotation forms a covariant approximate quantum error-correcting code with transversal U(1) logical gates. Generalizing previous work on “thermodynamic codes” to general local spin and different irreducible representations using only properties of the angular momentum algebra, we obtain bounds on the code inaccuracy under generic noise on any known $d$ sites, under independent and identically distributed noise, and under heralded $d$-local erasures. We demonstrate that this family of codes protects a probe state with quantum Fisher information surpassing the standard quantum limit when the sensing parameter couples to the generator of the U(1) logical gate.
@@
Motivated by these ideas, we introduce a family of AQECCs with transversal continuous gates, based on simple symmetry considerations. Additionally, these codes provide a metrological entanglement advantage for certain code parameters. More specifically, we consider NN spin-$s$ degrees of freedom and the irreducible representations (irreps) of their total SU(2) rotation. The irrep basis is $|J,\!M\rangle$, where $J\leq sN$ is the total spin quantum number, and $M$ is the magnetic quantum number. We show that the states $|J,\!M\rangle$, with certain choices of $M$, form an AQECC with transversal U(1) logical gates, and provide bounds on its code inaccuracy under generic noise on any known $d$ sites, under independently and identically distributed (i.i.d.) noise for $J\leq sN$, and under heralded $d$-local erasures for $J=sN$.
Our results stem from generalizing and improving upon results in Refs. [26, 20, 10, 25]. A special case of our AQECCs is $s=\frac{1}{2}$ and $J=sN=\frac{N}{2}$, in which case $|J\!,\!M\rangle$ are the Dicke states [27], and the corresponding code is the “thermodynamic code” [26, 20, 10, 25]. Its code-inaccuracy bound against generic noise on any known $d$ sites was first analyzed in Ref. [26], which also reported a coincidental identical bound for the $(s,J)=(1,N)$ code. We generalize this code to general $s$ and $J\sim N^{a}$ ($a\leq 1$) and tighten the bound against the same noise. We obtain our bound using the SU(2) ladder algebra without explicit expressions for the code words, therefore producing a general $s$- and $J$-dependent bound for this family of codes and explaining the coincidence from Ref. [26].
[…]
The inaccuracy bound of the “thermodynamic code” with code words $|M|=O(1)$ in the large-$N$ limit against heralded $d$-local erasure errors for a constant $d$ is reported in Refs. [20, 25]. We generalize it to $s\geq\frac{1}{2}$ and $J=sN$, but with $|M|=O(N^{b})$ and error locality $d\sim N^{c}$ for some exponents $b$ and $c$. We obtain a general $s$-dependent code-inaccuracy bound in this case using the Clebsch-Gordan coefficients without writing out the code words.
@@
Code words from SU(2) irrep states—Consider a system of NN spin-$s$ degrees of freedom (i.e. qudits of dimension 2s+12s\!+\!1), where $s$ can be a half integer or an integer. The total SU(2)-rotation unitary is a tensor product of the SU(2) rotations on the individual spins, generated by the su(2) algebra
@@
where $j$ labels the spin and $\hat{q}_{j}^{+},\hat{q}_{j}^{-}]=2\hat{q}^{z}_{j}$, $\hat{q}_{j}^{z},\hat{q}_{j}^{\pm}]=\pm\hat{q}_{j}^{\pm}$. This yields $\hat{Q}^{+},\hat{Q}^{-}]=2\hat{Q}^{z}$ and $\hat{Q}^{z},\hat{Q}^{\pm}]=\pm\hat{Q}^{\pm}$.
The irreps of the total SU(2)-rotation unitary are labeled by $J\leq sN$, with a possible multiplicity, and each irrep has dimension 2J+12J\!+1 with the basis labeled by $|J\!,\!M\rangle$, $M\in\{-J,\cdots,J\}$. The state $|J\!,\!M\rangle$ satisfies
@@
We will show that the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, equipped with transversal U(1) logical gates generated by $\hat{Q}^{z}$, forms an AQECC under generic noise on any known $d$ sites, i.i.d. error with sufficiently low error probability, and heralded $d$-local erasures for some parameters—provided the spacing $\Delta$ between magnetic quantum numbers is large enough. For the maximum $J=sN$, $|J\!,\!M\rangle$ is totally permutation symmetric, and the irrep multiplicity is one. For a general $J<sN$, one can have multiple irreps with the same $J$, and we consider $|J\!,\!M\rangle$ in any irrep or from any linear superposition of different irreps with the same $J$, as long as the ladder algebra in Eq. (2) holds among the resulting $|J\!,\!M\rangle$. As such, we suppress the multiplicity index.
Code inaccuracy against generic errors—We use the notation ((N,k))(\!(N,k)\!) and $\epsilon(\mathcal{N})$ for an AQECC, where $N$ denotes the number of qudits, $k$ denotes the number of logical qubits, and $\epsilon(\mathcal{N})$ denotes the optimal recovery inaccuracy against noise channel $\mathcal{N}$ (with encoding):
@@
where $\mathcal{R}$ is the recovery channel and $D_{P}(\mathcal{E},\mathcal{F})$ is the purified distance between channels $\mathcal{E}$ and $\mathcal{F}$, defined as
$D_{P}(\mathcal{E},\mathcal{F})=\max_{|\psi\rangle}\sqrt{1-f^{2}[\mathcal{E}\otimes\mathrm{id}(\psi),\mathcal{F}\otimes\mathrm{id}(\psi)]}~,$ (4)
where $f(\rho,\sigma):=\mathrm{Tr}[\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}]$ is the fidelity, $\psi:=|\psi\rangle\langle\psi|$, $|\psi\rangle$ is a purification of $\rho$, and $\mathrm{id}$ is the identity channel on the auxiliary space whose input has the same Hilbert space dimension as the input of $\mathcal{E}$ and $\mathcal{F}$ (see Supplemental Material S1 [40] for a list of definitions used throughout the paper). As we only consider noise on at most $d$ sites, we use $\epsilon(d)$ to stress its dependence on $d$, while suppressing $\mathcal{N}$ whose specific choices should be clear from the context.
@@
where $\lambda_{ij}$ are the components of a density operator, and $D_{P}(\Lambda+\mathcal{B},\Lambda)\leq\delta$, where $\Lambda(\rho)=\sum_{i,j}\lambda_{ij}\mathrm{Tr}(\rho)|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\Lambda(\rho)+\sum_{i,j}\mathrm{Tr}(\rho B_{ij})|i\rangle\langle j|$.
@@
Physically, this lemma holds since a $d$-local operator can change the magnetic quantum number $M$ by at most $2sd$. It ensures the “off diagonal” part of the Knill-Laflamme conditions to be exactly zero if the code words have enough spacing in $M$. The proof is simple and given in Supplemental Material S2 [40].
@@
Assume $\hat{F}$ is a $d$-local operator, then $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|\leq dq_{0}\|\hat{F}\|_{op}C(n,m)$, where $C(n,m):=\sum_{M=n}^{m-1}[c^{+}_{M}]^{-1}$ and $q_{0}$ is a constant. In particular, for $0\leq b<a\leq 1$, if we take $J\sim N^{a}$ and $|m-n|\sim N^{b}$, then we have $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|=\|\hat{F}\|_{op}\cdot O(dN^{-(a-b)})$.
Note that $\|\hat{F}\|_{\text{op}}:=\sup_{|\psi\rangle}\{\|\hat{F}|\psi\rangle\|:\||\psi\rangle\|\!=\!1\}$ is the operator norm and $\||\psi\rangle\|:=\sqrt{\langle\psi|\psi\rangle}$. Relegating the proof to Supplemental Material S3 [40], we comment that this bound is obtained by bounding $|\langle J,n+1|\hat{F}|J,n+1\rangle-\langle J,n|\hat{F}|J,n\rangle|$ and using the triangle inequality. We emphasize that the proof only uses the ladder algebra Eq. (2) and the locality of the operator $Q^{\pm}$ without the explicit expression for $|J\!,\!M\rangle$.
Using these two lemmas, we obtain the following upper bound on the code inaccuracy for generic noise on any known $d$ sites $\mathcal{N}(\rho)=\sum_{j=1}^{n_{d}}K_{j}(\rho)K^{\dagger}_{j}$ whose Kraus operators $K_{j}$ are at most $d$ local, acting on a fixed set of $d$ qudits, and $n_{d}=(2s+1)^{2d}$ is the number of independent Kraus operators. The proof is presented in Supplemental Material S4 [40] .
@@
Consider the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\!\sim\!N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\!\sim\!N^{b}$, and $b\!<\!a\!\leq 1$. The code $\mathfrak{C}$ forms an $(\!(N,O[b\log_{2}N])\!)$ AQECC against $d$-local noise on known sites, with inaccuracy $\epsilon(d)=O[\sqrt{d}(2s+1)^{d}N^{-(a-b)/2}]$ if $\Delta\geq 2sd+1$.
Figure 1: (a) For our SU(2) code with total spin $J\!\sim\!N^{a}$ and magnetic quantum number range $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, under heralded $d$-local noise channel ($d\!\sim\!c\log_{2}N$), the $(b,c)$ region where $\epsilon\!\to\!0$ (blue and red areas), and the region with metrological advantage (red area). (b) For $J=sN$, $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$, and $d$-local erasures ($d\sim N^{c}$), the $(b,c)$ region where $\epsilon\to 0$ (blue and red areas), and the region with metrological advantage (red area), all independent of $s$.
We focus on parameters where $\epsilon\rightarrow 0$ in the thermodynamic limit $N\rightarrow\infty$. For error size $d\!\sim\!c\log_{2}N$, we have $\epsilon=O(\sqrt{\log_{2}N}N^{-[a-b-2c\log_{2}(2s+1)]/2})$, which vanishes as $N\rightarrow\infty$ provided $c<(a-b)/[4\log_{2}(2s+1)]$. This is illustrated in Fig. 1(a) as the shaded area.
@@
The reason to consider the noise on a fixed set of $d$ sites in Theorem 1 is to control the number of Kraus operators $n_{d}$. For generic $d$-local noise $\mathcal{N}(\rho)=\sum_{j=1}^{n_{d}}K_{j}(\rho)K^{\dagger}_{j}$, where $K_{j}$ can act on different sets of $d$ qudits, $\mathfrak{C}$ forms an $(\!(N,O[b\log_{2}N])\!)$ AQECC with $\epsilon(d)=O(\sqrt{d}n_{d}N^{-(a-b)/2})$ if we take $\Delta\geq 4sd+1$. (See Supplemental Material [40] S4 for Corollary 1.)
Physically, the above result implies that the code can tolerate any error with the number of Kraus operators growing slower than $O(\sqrt{d}N^{-(a-b)/2})$, assuming each has operator norm $O(1)$. However, in practice, more nonlocal errors tend to occur with lower probability, resulting in smaller operator norms for their corresponding Kraus operators. This structure allows the code to tolerate noise channels with a larger number of Kraus operators. Indeed, under i.i.d. errors at unknown locations, we show in Supplemental Material [40] S5 (Corollary 2) that our code can achieve vanishing code inaccuracy by correcting $d$-local errors with $d\sim c\log_{2}N$, provided the error rate is sufficiently low $p\lesssim N^{-2}\log^{2}N$. Although this error rate implies a vanishing expected number of errors $Np\lesssim N^{-1}\log^{2}N$, it is chosen to enable a rigorous analytical bound on the code inaccuracy. It remains an interesting open question whether the code can still achieve low inaccuracy when the number of errors becomes nonvanishing.
Code inaccuracy against heralded erasures—The inaccuracy of the “thermodynamic code” against heralded $d$-local erasures for $d=O(1)$ and $|M|=O(1)$ was found in Refs. [20, 25] to be $\epsilon=O(N^{-1})$. The heralded $d$-local erasure channel is $\mathcal{N}_{\alpha}(\rho):=\mathrm{Tr}_{\alpha}[\rho]$, where $\alpha$ denotes the set of $d$ erased qudits.
We now generalize the code-inaccuracy bound against heralded erasures to $s$, $d\sim N^{c}$ and $|M|=O(N^{b})$. In this case, we consider codes with $J=sN$ only, where $|J\!,\!M\rangle$ is totally permutation symmetric. Owing to the irrep multiplicity being one, the $J=sN$ assumption allows us to express each codeword uniquely in terms of a basis for two subsystems using Clebsch–Gordan coefficients. To bound the code inaccuracy, we follow Ref. [18] by bounding the channel distance via the complementary channel $\hat{\mathcal{N}}_{\alpha}(\bullet)=\mathrm{Tr}_{\bar{\alpha}}(\bullet)$, where $\bar{\alpha}$ denotes the set of qudits not in $\alpha$ (the complement of $\alpha$). This can be achieved by the following lemma.
@@
Consider the irrep of SU(2) formed by $|J\!,\!M\rangle$, where $J=sN$, and the reduced density matrix $\rho_{M}:=\mathrm{Tr}_{\bar{\alpha}}[|J\!,\!M\rangle\langle J\!,\!M|]$, where $\bar{\alpha}$ is the complement of the set of erased qudits $\alpha$. We have the fidelity $f(\rho_{M},\rho_{M=0})=1-O(\frac{dM^{2}}{sN^{2}})=1-O(N^{2b+c-2})$ asymptotically for $M\sim N^{b}$, $d\sim N^{c}$, and $1>b\geq c$.
Physically, the above lemma states that the environment can barely tell the code words apart from each other. In the proof (presented in Supplemental Material S6 [40]), we use Clebsch-Gordan coefficients without writing the code words out explicitly, enabling a general $s$-dependent bound on the fidelity for J=sNJ\!=\!sN. In particular, we can rewrite $|J\!,\!M\rangle=\sum_{m_{1}=-j_{1}}^{j_{1}}C^{JM}_{j_{1},m_{1},j_{2},m_{2}}|j_{1}\!,\!m_{1}\rangle_{\alpha}\otimes|j_{2}\!,\!m_{2}\rangle_{\bar{\alpha}}$, where $m_{2}=M-m_{1}$, $j_{1}=sd$, $j_{2}=J-j_{1}$, and $C^{JM}_{j_{1},m_{1},j_{2},m_{2}}=(\langle j_{1},m_{1}|\langle j_{2},m_{2}|)|J,M\rangle$ is the Clebsch-Gordan coefficient. The fidelity between $\rho_{M}$ and $\rho_{M=0}$ can therefore be calculated accordingly.
@@
Consider the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J=sN$, $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $b<1$. $\mathfrak{C}$ forms an $(\!(N,O[(b\!-\!c)\log_{2}N])\!)$ AQECC against heralded $d$-local erasures of size $d\sim N^{c}$, where $0\leq c\leq b<1$, with $\epsilon(d)=O(N^{-(1-b-c/2)})$ if we take $\Delta\geq 2sd+1$.
Theorem 2 can be obtained by combining our Lemma 1, our Lemma 3, and Theorem 3 of Ref. [20]. To make our presentation self-contained, we present the proof of Theorem 2 tailored to our case in Supplemental Material S7 [40]. References [20, 25] showed that $\epsilon=\Omega(N^{-1})$ for any U(1) covariant code. Thus, this family of SU(2) AQECCs is asymptotically optimal for general $s$ since $\epsilon=O(N^{-1})$ when $b=c=0$. Although Theorem 1 also applies to heralded erasure errors, the improvement in Theorem 2 arises from the assumption that $J=sN$, enabling a stronger result.
For general $b$ and $c$, to have $\epsilon\rightarrow 0$ when $N\rightarrow\infty$, we need c<2(1b)c<2(1-b). Additionally, we need $\Delta=\Omega(N^{c})$ for Lemma 1 and $c\leq b$ to ensure at least one logical qubit. This parameter regime is shaded in Fig. 1(b), which is independent of $s$ and $k$. The code words with $M\sim N^{b}$ for $b>\frac{1}{2}$ [red shaded area in Fig. 1(b)] are crucial for a metrological entanglement advantage.
Metrological entanglement advantage—We now apply our codes to quantum metrology and show that they can have an entanglement advantage with noise resistance. In the local parameter estimation problem [10], we would like to estimate the parameter $\theta$ imprinted on the probe state $|\psi_{0}\rangle\in\mathfrak{C}$ as $|\psi_{\theta}\rangle=e^{-i\hat{Q}^{z}\theta}|\psi_{0}\rangle\in\mathfrak{C}$. Abbreviating $\psi_{\theta}:=|\psi_{\theta}\rangle\langle\psi_{\theta}|$, the noisy probe state is then $\rho_{\theta}:=\mathcal{N}(\psi_{\theta})$ for the noise channel $\mathcal{N}$.
In quantum metrology, QFI is a central figure of merit, which characterizes how well a parameter can be estimated via the quantum Cramér-Rao bound [42]. For the states $\rho_{\theta}$ parametrized by $\theta$, the QFI is defined as
$\mathfrak{F}(\rho_{\theta};\partial_{\theta}\rho_{\theta}):=\mathrm{Tr}(\rho_{\theta}R^{2})~,$ (6)